With a laser, fringes remain clearly visible, when the path difference was increased up to X m, Deduce Coherence time and spectral half width given λ = 11.5x10-7 m.
X = 1.768 T = 77323
1
Expert's answer
2020-05-21T13:11:09-0400

As per the question,

Wavelength of the laser beam "(\lambda)=11.5\times 10^{-7}m"

Speed of light "(c)=3\times 10^8 m/sec"

"l_{coh}=1.768m"

"t_{coh}=\dfrac{l_{coh}}{c}=\dfrac{1.768}{3\times 10^{8}}sec=5.9\times 10^{-9}=5.9ns"


"\Delta \lambda=\dfrac{\lambda^2}{l}=\dfrac{(11.5\times 10^{-7})^2}{1.768}=74.8\times 10^{-14}m=7.48\times 10^{-15}m"